SI3 STATUS AT PURDUE Oct 27 1999
STATUS OF CLAMSHELL #1
STATUS OF CLAMSHELL #2
CLAMSHELL #2 IS ON HOLD PENDING REPLACEMENT PARTS SINCE 9/15/99
Replacement part, FL213 has had diamond added 10/26 tested fine it is ready for mounting (note added after meeting: part was mounted and tested fine 10/28: CLAMSHELL IS COMPLETE!)
STATUS OF LAYER 3
STATUS OF L3 BY LADDER
STATUS OF LAYER 4
To complete L4 we need to build:
5HL need to be built, all wafer chains for these HLıs are in process but starters not yet at Purdue
COMMENTS ON LAYER 4 HALF LADDERS
HL539 after p-side wire-bonding noisy channels, 47,12 n, p with n/p correspondence. Currents are normal. Can live with this if CU is happy. HL534 high currents (60 micro n/p) due to crater bond-pad. Status: use as is
HL526 high current (20 micro n/p) no optical inspection yet status: use as is
HL512 special long p-strip wire status OK
HL559 cracked and repaired hybrid status OK
HL553 used extensively for vee testing status OK
HL554 fried at Purdue status OK
HL506 not able to readout at Purdue believed to be due to incompatibility with system status: use
STATUS OF LAYER 4 BY LADDER